OpenStudy (twizttiez):

What is the y-intercept of the line? y-7x+1 y-intercept=____?

OpenStudy (anonymous):

1

OpenStudy (twizttiez):

Are you sure?

OpenStudy (anonymous):

wait is the equation y=-7x+1

OpenStudy (anonymous):

cause the = is important

OpenStudy (twizttiez):

Yes jdizzle

OpenStudy (twizttiez):

I miss typed sorry

OpenStudy (anonymous):

then I am positive it is 1

OpenStudy (anonymous):

and it is np

OpenStudy (twizttiez):

Thanks

OpenStudy (hyper):

To find the y-intercept, just substitute x = 0 into the equation...

OpenStudy (anonymous):

yup need anything else?

OpenStudy (anonymous):

 or this is always the y intercept

OpenStudy (anonymous):

in point slope form

OpenStudy (texaschic101):

JDizzle is correct...y int = 1

OpenStudy (twizttiez):

Ok thanks can you help me with some more?

OpenStudy (texaschic101):

what ya got ?

OpenStudy (twizttiez):

What is the slope, and what is the y-intercept of the line? y=2/3x-4 ? slope= ------ ? y-intercept=_____?

OpenStudy (anonymous):

slope=2/3 y-int=-4

OpenStudy (texaschic101):

In y = mx + b form, the slope will be in the m position and the y intercept will be in the b position. y = mx + b y = 2/3x - 4 slope = 2/3 y intercept = -4

OpenStudy (twizttiez):

Ok thanks ^_^ i hate to ask but i only need help on two more any way you guys could stay and help?

OpenStudy (texaschic101):

ask away

OpenStudy (anonymous):

im down :)